Given f(x)=3x+2 and g(x)=2x squared +1. Find f[g(2)]. Thank you for your help.
Answer by math-vortex(648)

Hi, there--

The Problem:
Given f(x)=3x+2 and g(x)=2x^2 +1. Find f[g(2)]. 

Solution:
f(x) = 3x+2
g(x) = 2x^2+1

We are going to work from the inside out. First, find g(2).
g(2) = 2(2)^2+1 = 2(4)+1 = 8+1
g(2) = 9

When x=2, the value of g is 9. Now we input 9 into the function f. 
f[g(2)] = f(9)
f(9) = 3(9)+2 = 27+2
f(9) = 29

f[g(2)] = 29


If this is unclear, or if you still have questions, email me.

Mrs.Figgy
math.in.the.vortex@gmail.com